### v1

@@ 4.3 Roles, Agents, and Services
Figure 5: Agents and their primary services.
From the scenario described in Section 4.1 and its subsequent refinement processes (see Figures 3 and 4), we identified five distinct types of roles: suppliers, wholesalers, retailers, logistics providers, and 3PL provider, and each with their corresponding primary services, which are a coherent blocks of functions that roles are expected to accomplish. Each of these roles can be represented by individual agents, serving as the structural entities within the scope of their respective organisations. These agents and their main services are shown in Figure 5. A detailed description of each agent are as follows:
@@
    Wholesaler Agent: This agent represent the wholesaler, which is the central roles, namely the CMC, in this scenario. Its main services include managing the processes of procuring meat from supplier agents and wholesaling meat to retailer agents, and corresponding inventory management during these two functions. It also interacts with logistics agents to arrange delivery service for its customers. This agent serves as a central hub connecting upstream and downstream supply chain agents, facilitating the entire flows.
@@
    Supplier Agent: This agent is responsible for managing the process of supplying meat to wholesaler agents on behalf of a supplier. Similar to wholesaler agents, the supplier agent provides its customers with delivery service, which is outsourced to logistics agents. Therefore, this agent needs to coordinate with wholesale agents and logistics agents to fulfil purchase orders.
@@
    Retailer Agent: Representing local retailers, such as restaurants or local stores, this agent purchases meat from wholesaler agents. It interacts with the wholesaler agent to handle procurement on behalf of the retailer.
@@
    Logistics Agent: This agent, representing a logistics company, provides logistics services to supplier agents and wholesaler agents in the scenario. Logistics companies primarily manage overall logistics service but outsource fulfilment service to 3PL providers. Logistics agents need to coordinate with suppliers, wholesalers, and 3PL providers for delivery arrangements.
@@
    3PL Agent: This agent, acting on behalf of a 3PL provider, is responsible for providing delivery services to logistics companies, fulfilling delivery orders assigned by a logistics agent. It monitors the entire transportation process of the meat, providing logistics companies with real-time data on the truck’s geolocation and the meat’s ambient conditions for further use, such as fulfilment service evaluation.
In addition to the main services discussed above (indicated by solid-lined ellipses in Figure 5), these agents also offer information query services to other agents and include functions for (un)registration. Information query service provide these agents with the capability to connect with other agents to request specific information. Moreover, they have the ability to register their services into the system, making them ready for discovery and use. Conversely, they can unregister existing services, removing them from the running system and making them unavailable for discovery by other agents.
These five types of agents are the main stakeholders — structural entities — in this meat supply chain, and their interactions shape the dynamics of this supply chain. Additionally, we introduced an functional agent, known as the admin agent, responsible for managing agent-level administration services, including providing yellow page services. The following sections provide a more detailed description of the organisation and interactions among these agents.
@@ 4.4 Connection Mechanism
Figure 6: Interaction diagrams of the contract net protocol and the HTTP protocol. Note: The left is the FIPA contract net interaction protocol, which is a slight adaption of the original contract net protocol [Smith and Davis, 1981].
Before distributed agents can be organised together, the connections between them must be first enabled. We, thus, first introduce connection mechanisms for facilitating agent organisation and interactions. Agents in an MAS offer services to others or request services from others. Therefore, this leads to the challenge of how a service requester can find suitable service providers capable of executing a service. This challenge is known as the connection problem [Davis and Smith, 1983, Ameri and McArthur, 2013].
In this prototype, we devised two mechanisms for agent connection. The first is the Open Economic Framework (OEF)-based connection, in which two agents are linked through a third-party service known as the OEF. In this type of connection, a service requester submits a service query to the OEF, which responds by providing a list of potential service providers that match the query criteria. The service requester can then choose one ore more service providers from this list to establish further connections. The OEF-based connection is an indirect or mediated connection, providing a means for two unknown agents to discover each other through the assistance of a mediator. The second mechanism is the HTTP-based connection, in which two agents connect directly through the HTTP protocol, as explained in Section 4.6. In this approach, a service requester sends an HTTP request to a service provider via its URL and receives an HTTP response from that service provider. This represents a direct connection, which is well-suited for scenarios in which both agents are already known each other and can contact each other directly without the need for an intermediary.
@@ 4.5 Agent Organisation
Connection mechanisms enable agents to establish connections with others. To achieve coherence, connections among agents should be organised as certain topology. As described in Section 4.3, this prototype includes five types of structural agents and an functional agent. In designing its agent organisation, we have chosen not to include intermediary agents, such as blackboard, yellow page, and broker, for two primary reasons: a) this prototype involves a small number of agents, and b) the OEF is capable of providing efficient discovery and matchmaking services for agents. The resulting agent organisation is generally a collective of peers, as illustrated in Figure 7. These peers operate with independent control regimes, each focusing on specific stages of the three flows. The two connection mechanisms discussed in the previous section are used to connect these peers, denoted by solid and dashed lines, respectively. The admin agent can establish connections with all other agents via the HTTP connection, thereby maintaining an portal that provides an overview of the status of all agents in the system. The supplier, wholesaler, and retailer agents are linked by the OEF connection, enabling negotiations for replenishment or procurement. The HTTP connection becomes active between these agents if trust is established between them.
@@ 4.6 Agent Interactions, Language and Protocols
While the agents and their organisation define the structure of the A2SC system, this section discusses its dynamics, involving the interactions among these agents and the language and protocols that facilitate these interactions. As shown in Figure 4, the wholesale agent interacts with supplier agents in the replenishment process and with retailers during the wholesale process. Both processes require logistics services, thus involving two additional agents: logistics and 3PL provider agents. These agents interact by adhering to agreed-upon interaction protocols.
@@
In this implementation, we incorporate two types of interaction protocols for facilitating interactions: the contract net protocol and the HTTP protocol. The contract net protocol is specifically designed for scenarios where a service requester needs to select suitable service providers capable of performing a specific task. This protocol often include multiple rounds of negotiation and information exchange, as illustrated in Figure 5(a). It is well-suited for managing complex interaction scenarios that may involve unknown participants, such as procurement processes. On the other hand, the HTTP protocol is employed for simpler interaction scenarios that typically involve require only a single round of interaction. As shown in Figure 5(b), a service requester, referred to as an HTTP client, can send a request to a service provider, an HTTP server, to either get data or post a task for execution. This protocol is most suitable for interactions between known participants who have already established a level of trust. For example, a logistics agent may directly assign a delivery task to one of its longstanding 3PL partners for fulfilment. Additionally, the HTTP protocol is appropriate for quickly retrieving information such as product wholesale price, delivery rates, and current traffic conditions.
@@
The interactions among agent in the two processes are depicted in Figures 8 and 9, In both figures, both processes involve four types of agents, and these agents interact with each other through messages. The messages in the two figures are composed of two essential parts: performative and message body, delimited by a colon. The two parts respectively represent the communicative act (behaviour) and the content of the message. Interaction protocols define a set of performatives that may be used in the interaction. For instance, the contract net protocol includes performatives such as cfp (call for proposal), propose, and reject-proposal, while the HTTP protocol includes two basic performatives: request (which further has two types of requests: get and post) and response, as shown in Figures 8 and 9. Additionally, a message should include other metadata for unambiguous communication, such as sender, receiver, protocol, and ontology. The agent communication language defines the message format, and in this study, we we use FIPA-ACL as the language for agent communication.
@@
Both processes involves multiple rounds of communication, sending and receiving messages to coordinate procurement and logistics service. They also include procedures to support decision-making, such as assess_order() for evaluating incoming procurement orders, and self-calls for executing specific tasks, such as monitor_delivery() for tracking delivery process. In the replenishment process, which offers multiple delivery options, there are additional rounds of interaction with the logistics agent, and subsequently with the 3PL agent, to request information about available delivery options. This is depicted in the right bottom area in Figures 8. Although the two processes are operationally separate and independent, the wholesale process may trigger the replenishment process automatically. For instance, when the wholesaler’s inventory reaches a predetermined low stock level, the replenishment process is initiated. In this sense, the two processes are interconnected, illustrating an end-to-end supply chain that facilitates the automated movement of “goods” from suppliers through intermediaries to customers.
@@ 4.7 Implementation
The previous sections provide an overview of the A2SC prototype’s design. In this section, we will first introduce the toolkit for developing this prototype, and then delve into its implementations.
@@ 4.7.1 Toolkit for Implementation
We used Python as the primary language for developing the backend of this prototype, while other languages such as HTML and JavaScript were used for creating the frontend interfaces. To streamline the development process, we leveraged a toolkit consisting of the following tools and frameworks:
  • AEA Framework: The AEA (Autonomous Economic Agent) framework [Minarsch et al., 2021] from Fetch.ai  [^5] https://fetch.ai/ technology stack is a Python-based toolkit for constructing autonomous agents known as AEAs. These AEAs operate independently on behalf of their owners and autonomously execute actions to achieve predefined objectives. We used this framework to develop the agent network in the prototype.
  • OEF (Open Economic Framework): The OEF serves as a decentralised communication, search and discovery system for AEAs, providing protocols, languages, connection mechanisms that facilitate agents search, interaction discovery, and engagement within established social norms. It establishes a virtual environment where the inhabitants, AEAs, can operate and interact. In Fetch.ai technology stack, the AEA framework and the OEF represent the top two layers, positioned above the smart ledger layer, as depicted in Figures LABEL:fig:fetch_three_layers. For this implementation, we exclusively utilised the AEA framework and the OEF to develop the prototype, as they are capable of functioning independently without the smart ledger.
  • REST and OpenAPI: REST (REpresentational State Transfer) Fielding [2000] is an architectural style for providing interface standards between decoupled components on the web. RESTful APIs are typically based on HTTP methods to access resources via URL-encoded parameters and the use of JSON or XML for data transmission. All agents in the prototype were equipped with RESTful APIs, enabling them communicate with each other without the OEF and access external resources. However, these agents need to expose their APIs to other in a standard manner. The OpenAPI specification (OAS) [OpenAPI Initiative, 2021] was therefore adopted for describing these interfaces. The OAS defines a standard, programming language-agnostic interface description for RESTful APIs, allowing both humans and machines to discover and comprehend a web service without requiring access to source code and extra documentation. In this prototype, we define interface description files (.yaml) by adhering to the OAS.
  • WebSockets: WebSockets, which enable bi-directional, full-duplex communication, are well-suited for frequently updated or real-time applications. We thus used WebSockets [Fette and Melnikov, 2011] for services such as notification, monitoring, and agent chatting. The above tools are relevant to backend development and architectural design. We will describe the tools/frameworks used for frontend development in this following paragraphs.
In addition to these major tools, we incorporated complementary tools and frameworks such as Bootstrap  [^11] https://getbootstrap.com/, JQuery  [^12] https://jquery.com/, and Ajax. This toolkit instrumental in enhancing communication, interaction, and visualisation functionalities for both backend agent network and frontend interfaces during the development of the A2SC prototype.
@@ 4.7.2 System Architecture
The A2SC prototype primarily consists of two main parts: an agent network and an interface web application. We implemented a microservice architecture to integrate these components using RESTful APIs and WebSockets, as depicted in Figure 10. The agent network is composed of a collection of AEAs, with each representing distinct organisations or functions. These AEAs operate within an “environment”, specifically the OEF, and collaborate cohesively to pursue their respective objectives. Additionally, they may also access storage services to persist data and utilise sensors to sense the environment. The interface application, powered by Django and Channels, provides a web-based portal to the prototype. The interface comprises a client with a collection of UI components, which we will further described in the next section. Additionally, it includes a server that handles HTTP and WebSockets requests through its Views and Consumers modules, respectively. Specifically, it handles user input and provides real-time visualisation of the agent network’s state through RESTful APIs and/or the WebSocket protocol. This architectural design enables a loose coupling between the agent network, which may be distributed over many machines, and the interfaces, allowing them to maintain independence and extensibility.
@@ 4.7.3 Interface
Based on the predefined scenario and the system architecture, specifically the left part presented in Figure 10, we developed a single-page interface for this prototype. The web-based interface is constructed using a Channels-extended Django framework, enabling access through a web browser. The interface consists of four main panels, each serving a specific purpose:
@@
    Logistics Monitoring: This panel visualise the real-time track and trace of the vehicle fulfilling the delivery.
@@
These panels interact with the agent network using either HTTP or WebSockets. We use Ajax for the ordering panel and WebSockets for the other other panels to dynamically update the current interface with new data from the backend systems, enhancing use experience by eliminating the need to load entire new pages.
@@ 4.7.4 Data Preparation
As previously described, this prototype involves real-time data collection to monitor meat delivery processes. Instruments such as sensors and GPS trackers are required to capture the ambient condition of the meat and the location of its transportation vehicle. This data needs to be continually collected during the delivery process. However, due to limited experimental condition, it is not feasible to actually drive a vehicle with necessary data collection instruments and refrigerated containers to transport meat from one place to another while establishing a wireless communication link with remote agents. Moreover, using a realistic setting for this prototype would be uneconomical and unnecessary, particularly in light of the primary objective of this study, which is to showcase the technical methodology rather than achieving a high level of technical readiness.
To address this, we used pre-collected IoT and GPS data to simulate real-time data collection during the delivery process. These data were collected through the following steps:
@@
    Install a UbiBot all-in-one sensor  [^13] https://www.ubibot.com/ubibot-ws1/ and a smartphone with a GPS recording App (myTracks [^14] https://apps.apple.com/us/app/mytracks-the-gps-logger/id358697908) on a bike.
@@
The collected data consists of a series of IoT (temperature, humidity, and light) and geolocation data (longitude, latitude, and elevation) points indexed in chronological order, essentially forming time series data. We utilised these CSV files to simulate real-time data capture during the delivery monitoring.
@@ 4.7.5 Implementation Details
Following the design described previously, we implemented this prototype using Python and the toolkit presented in Section 4.7.1. In this section, we will delve into the specifics of the implementation.

### v2

@@ 4.3 Roles, Agents, and Services
Figure 5: Agents and heir main
services.
From the scenario described in Section  4.1 and its subsequent refinement processes (see Figures 3 and 4), we identified five distinct types of roles: suppliers, wholesalers, retailers, logistics providers, and 3PL provider, and each with their corresponding primary services, which are a coherent blocks of functions that roles are expected to accomplish. Each of these roles can be represented by individual agents, serving as the structural entities within the scope of their respective organisations. These agents and their main services are shown in Figure 5. Each agent is detailed as follows:
@@
    Wholesaler Agent: This agent represent the central role of the wholesaler (the CMC) in this scenario. Its main services involve managing the procurement of meat from supplier agents, wholesaling meat to retailer agents, and overseeing inventory during these processes. It also coordinates with logistics agents to arrange delivery service for customers, functioning as a central hub connecting upstream and downstream supply chain agents, facilitating entire flows.
@@
    Supplier Agent: This agent supplies meat to wholesaler agents on behalf of a supplier. Like wholesaler agents, the supplier agent outsources delivery services to logistics agents, necessitating coordination with both wholesale agents and logistics agents to fulfill purchase orders.
@@
    Retailer Agent: Representing local retailers, such as restaurants or local stores, this agent procures meat from wholesaler agents and interacts with them to manage procurement on behalf of the retailer.
@@
    Logistics Agent: This agent, representing a logistics company, offers logistics services to supplier agents and wholesaler agents in the scenario. Logistics companies manage overall logistics service but outsource fulfilment service to 3PL providers. Logistics agents coordinate with suppliers, wholesalers, and 3PL providers for delivery arrangements.
@@
    3PL Agent: Acting on behalf of a 3PL provider, this agent provides delivery services to logistics companies, fulfilling delivery orders assigned by a logistics agent. It monitors the entire meat transportation process, supplying real-time data on the truck’s geolocation and ambient conditions for logistics companies to use, such as in fulfilment service evaluation.
In addition to the main services discussed above (represented by solid-lined ellipses in Figure 5), these agents provide information query services to other agents and include functions for (un)registration. Information query service enable agents to request specific information from other agents. They can register their services for discovery or unregister existing services, removing them from the system and making them unavailable for discovery by other agents.
These five types of agents are the main stakeholders — structural entities — in this meat supply chain, and their interactions form the dynamics of this supply chain. Additionally, we introduced a functional agent, the admin agent, responsible for managing agent-level administration services, including providing yellow page services. The following sections describe the organisation and interactions among these agents.
@@ 4.4 Connection Mechanism
Figure 6: Interaction diagrams for the contract net protocol and the HTTP protocol. Note: The left is the FIPA contract net interaction protocol, which is a slight adaption of the original contract net protocol (Smith and Davis, 1981).
Prior to organising distributed agents, it is crucial to enable connections among them. We first introduce connection mechanisms to facilitate agent organisation and interactions. In an MAS, agents offer services to others or request services from others, matching service requester with suitable service providers. This is known as the connection problem (Davis and Smith, 1983; Ameri and McArthur, 2013) in distributed problem solving.
In this prototype, we devised two agent connection mechanisms. The first is the Open Economic Framework (OEF)-based connection, where two agents are linked through a third-party service, the OEF. In this indirect or mediated connection, a service requester queries the OEF, which responds with a list of potential service providers. The requester can then choose one or more providers from the response for further connections. The second mechanism is the HTTP-based connection, where two agents connect directly through the HTTP protocol (see Section  4.6). In this connection, a service requester sends an HTTP request to a service provider via its URL and receives an HTTP response. This connection is suitable for scenarios where both agents are already acquainted and can contact each other directly without an intermediary.
@@ 4.5 Agent Organisation
Connection mechanisms facilitate agent connections, and for coherence, these agents should be organised in a specific structure — agent organisation — which defines the way of agent communications and connections (Horling and Lesser, 2004; Dorri et al., 2018). As described in Section  4.3, this prototype includes five types of structural agents and a functional agent. These functional agents are self-interested but may collaborate for a short-term goal such as order fulfilment or build a long-term partnership for outsourcing logistics services. Intermediary agents such as blackboard, yellow page, and broker were not included in this design due to: a) this prototype involves a small number of agents, and b) the OEF is capable of providing efficient discovery and matchmaking services. Therefore, we employ a compound organisation paradigm (see (Horling and Lesser, 2004) for more details about this organisation), in which multiple organisation forms (such as coalition and team) are adopted for organising the relationship among agents, as illustrated in Figure 7. For better illustration, although unnecessary for an organisational diagram, we differentiate the two connections discussed earlier between agents in this diagram, denoting them with solid and dashed lines, respectively. These agents operate with independent control regimes, each focusing on specific stages of the three flows. The admin agent can establish connections with all other agents via the HTTP connection, maintaining an portal that provides an overview of the status of all agents in the system. The OEF connection links the supplier, wholesaler, and retailer agents, enabling negotiations for replenishment or procurement. The HTTP connection becomes active between these agents if trust is established.
@@ 4.6 Agent Interactions, Language and Protocols
While agents and their organisation define the A2SC system’s structure, this section discusses its dynamics, focusing on interactions among agents and the language and protocols facilitating these interactions. As shown in Figure 4, the wholesale agent interacts with supplier agents in the replenishment process and with retailers during the wholesale process. Both processes require logistics services, involving two additional agents: logistics and 3PL provider agents. These agents interact by following agreed-upon interaction protocols.
@@
In this implementation, we incorporate two types of interaction protocols: the contract net protocol and the HTTP protocol. The contract net protocol is specifically designed for scenarios where a service requester needs to select suitable service providers for executing a specific task, often involving multiple rounds of negotiation and information exchange, as illustrated in Figure 6(a). It is well-suited for managing complex interactions with unknown participants, such as procurement. Conversely, the HTTP protocol is adopted for simpler interactions typically involve requiring only a single round. As shown in Figure 6(b), an HTTP client (service requester) can send a request to an HTTP server (service provider) to either get data or post a task for execution. This protocol is most suitable for interactions between known participants with established trust, such as a logistics company assigning a delivery task to their longstanding 3PL partners. Additionally, the HTTP protocol is appropriate for quickly retrieving information such as product wholesale price, delivery rates, and current traffic conditions.
@@
The detailed interactions among agent in the two processes are illustrated in Figures 8 and 9, Both processes involve four types of agents, interacting through messages. Messages in these figures consist of two essential parts: performative and message body, separated by a colon. These parts respectively represent the communicative act (behaviour) and the content of the message. Interaction protocols define a set of performatives used in interactions. For example, the contract net protocol includes performatives such as cfp (call for proposal), propose, and reject-proposal. The HTTP protocol includes two basic performatives: request (further categorised as get and post) and response, as shown in Figures 8 and 9. A message should include metadata for unambiguous communication, such as sender, receiver, protocol, and ontology. The agent communication language defines the message format, and in this study, we use FIPA-ACL for agent communication.
@@
Both processes involves multiple rounds of communication, sending and receiving messages to coordinate procurement and logistics service. They also include decision-making support procedures, such as assess_order() for evaluating incoming procurement orders, and self-calls for executing specific tasks, such as monitor_delivery() for tracking delivery process. In the replenishment process, which offers multiple delivery options, there are additional rounds of interaction with the logistics agent, and subsequently with the 3PL agent, to request information about available delivery options. This is depicted in the right bottom area in Figures 8. Although the two processes are operationally separate and independent, the wholesale process may trigger the replenishment process automatically when conditions are met. For instance, when the wholesaler’s inventory reaches a predetermined low stock level, the replenishment process is initiated. In this sense, the two processes are interconnected, illustrating integrated supply chain processes that facilitate the automated movement of “goods” from suppliers through intermediaries to customers.
@@ 4.7 Implementation
The previous sections provide an overview of the A2SC prototype’s design. This section describes the implementation details of the prototype.
@@ 4.7.1 Toolkit for Implementation
We used Python as the primary language for developing the backend of this prototype, while other languages such as HTML and JavaScript were used for creating frontend interfaces. To enhance development efficiency, we employed a set of tools and frameworks detailed in A), facilitating communication, interaction, and visualisation functionalities for both the backend agent network and frontend interfaces in the A2SC prototype.
@@ 4.7.2 System Architecture
This A2SC prototype primarily consists of two main components: an agent network and an interface web application. We implemented a microservice architecture to integrate these components using RESTful APIs and WebSockets, as depicted in Figure 10. The agent network is composed of a collection of autonomous economical agents (AEAs), implemented using the Fetch AEA framework as described in A, with each representing distinct organisations or functions. These AEAs operate within an “environment”, specifically the OEF, and collaborate cohesively to pursue their respective objectives. Additionally, they may also access storage services to persist data and utilise sensors to sense the environment. The interface application, powered by Django and Channels, provides a web-based portal to the prototype. The interface comprises a client with a collection of UI components, which we will further described in the next section. Additionally, it includes a server that handles HTTP and WebSockets requests through its Views and Consumers modules, respectively. Specifically, it handles user input and provides real-time visualisation of the agent network’s state through RESTful APIs and/or the WebSocket protocol. This architectural design enables a loose coupling between the agent network, which may be distributed over many machines, and the interfaces, allowing them to maintain independence and extensibility.
@@ 4.7.3 Interface
We developed a single-page web-based interface for this prototype, based on the predefined scenario and the system architecture. The interface consists of four main panels, each serving a specific purpose:
@@
    Logistics Monitoring: This panel visualises the real-time track and trace of the vehicle fulfilling the delivery.
@@
These panels interact with the agent network using either HTTP or WebSockets. We use Ajax for the ordering panel and WebSockets for the other panels to dynamically update the current interface with new data from the backend systems, enhancing use experience by eliminating the need to load entire new pages.
@@ 4.7.4 Data Preparation
As previously described, this prototype involves real-time data collection to monitor meat delivery processes. Instruments such as sensors and GPS trackers are required to capture the ambient condition of the meat and the location of its transportation vehicle. This data needs to be continually collected during the delivery process. However, due to limited experimental condition, it is not feasible to actually drive a vehicle with necessary data collection instruments and refrigerated containers to transport meat from one place to another while establishing a wireless communication link with remote agents. The primary objective of this study is to demonstrate the technical methodology rather than achieve a high level of technical readiness. Employing a realistic setting for this prototype would be uneconomical and unnecessary.
We therefore used pre-collected IoT and GPS data to simulate real-time data collection during the delivery process. These data were collected through the following steps:
@@
    Install an all-in-one sensor  [^5] UbiBot WS1 was used in this experiment. and a smartphone with a GPS tracking App (myTracks  [^6] https://apps.apple.com/us/app/mytracks-the-gps-logger/id358697908) on a bike.
@@
The collected data consists of a series of sensor (temperature, humidity, and light) and geolocation (longitude, latitude, and elevation) data points indexed in chronological order. We utilised these data to simulate real-time data capture during the delivery monitoring.